500 tickets were sold for a certain performance of a play. The tickets for adults and children sold for $7.50 and $4.00, respectively, and the total receipts for the performance were $3312.50. How many of each kind of ticket were sold?
Answer by rfer(16322)

a+c=500 c=500-a
7.5a+4c=3312.50
7.5a+4(500-a)=3312.50
7.5a=2000-4a=3312.50
3.5a=1312.50
a=375
c=500-375=125